OpenStudy (anonymous):

Hey Ill give medals and fans if you help :D A new law requires that 12% of an individual's income be invested in the stock market. Your accounts show that you need to put $420 in the stock market this year. How much did you earn this year.

OpenStudy (anonymous):

470.40

OpenStudy (anonymous):

3500 because, 420/12 = 35 so, 35 * 100 = 3500 therefore the answer is 3500 so, if we were to do the opposite. 3500/10 = 35 35*12 = 420